# Implementing Cumulative Functions with Generalized Cumulative Constraints

### v1

@@ 2 $p_{i}\leftarrow true$
Another optimization is to avoid considering some tasks when performing the timetabling algorithm. Indeed, fixed tasks that occur before the start of the earliest unfixed task or after the end of the latest unfixed task are not useful for filtering. Thus, such tasks do not need to be processed which reduces the number of time points in the profile. To do so, we use the same process as the Fruitless Fixed Tasks Removal of (Gay, Hartert, and Schaus 2015a).
@@
  1. 1.
    When processing task $A$ at time point 111, its maximum height $\overline{c}_{A}$ is adjusted to 111 by the Height rule.
  2. 2.
    When processing task $B$ at time point 222, the task is detected as non-feasible at this time point by the Forbid rule. Its minimum start $\underline{s}_{B}$ is thus pushed to the next time point at time 333 where the adjustment stops as the task is feasible at this time. This also updates its minimum end to $\underline{e}_{B}=6$ due to the internal relation $\underline{e}_{B}\geq\underline{s}_{B}+\underline{d}_{B}$.
  3. 3.
    When processing task $C$ at time point 444, the task is detected as mandatory. The following adjustments are done: 1) The task is set to present. 2) Its maximum starting time is set to $\overline{s}_{C}=4$ which also updates its maximum ending time to $\overline{e}_{C}=7$ due to the internal relation $\overline{e}_{C}\leq\overline{s}_{C}+\overline{d}_{C}$. 3) Its minimum ending time is set to $\underline{e}_{C}=5$. 4) As the value $overload=-1$ is negative, the maximum height of the task is set to $\overline{c}_{C}=-1$. Note that even if task $B$ is processed before task $C$ and its fixed part increased, task $C$ is adjusted based on the profile range which has been computed based on the state of task $B$ at the start of the propagation.
@@
$A$
: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=1,\allowbreak p=true\rangle$
$B$
: $\langle s=[3,4],\allowbreak d=[3,4],\allowbreak e=[6,7],\allowbreak c=2,\allowbreak p=true\rangle$
$C$
: $\langle s=[3,4],\allowbreak d=[1,3],\allowbreak e=[5,7],\allowbreak c=[-2,-1],\allowbreak p=true\rangle$
Related Work
In (Beldiceanu and Carlsson 2002), a timetabling filtering for the GeneralizedCumulative constraint with alternative resources was introduced. It collects a list of tasks at relevant time-point before filtering them. Our method is more in line with the strategy described in (Gay, Hartert, and Schaus 2015a) for the classical cumulative, which also constructs a profile and then confronts each task with it. We instead use the Profile data structure (Gingras and Quimper 2016), to prune the tasks forward and backward. This reduces the number of calls required to reach the fix-point of the constraint. Contrary to (Gay, Hartert, and Schaus 2015a), which uses just a list of rectangles for the profile, our algorithm finds the starting rectangle in $\mathcal{O}(1)$ thanks to the Profile. Consequently, despite a worst-case quadratic time complexity per call similar to that described in (Beldiceanu and Carlsson 2002) and (Gay, Hartert, and Schaus 2015a), the proposed algorithm can be faster. Furthermore, our algorithm also prunes the height and the maximum length of tasks. More details on the filtering differences with the other implementations from the literature are provided in the technical Appendix.
Experimental Results
@@
Our approach and filtering algorithm was implemented in Java on MaxiCP, an extension of the MiniCP (Michel, Schaus, and Van Hentenryck 2021) solver. We compare it with the closed-source state-of-the-art commercial solver CPOptimizer (Laborie et al. 2018) as well as the open-source solver Gecode (Schulte, Lagerkvist, and Tack 2006) that implements a version of the GeneralizedCumulative constraint of (Beldiceanu and Carlsson 2002). As cumulative functions are not available in Gecode, we carefully modeled them in a way similar to how they would be flattened by Algorithm 1. Three cumulative problems that involve reservoirs or negative consumptions are considered. We use the same fixed static search strategy for all solvers and problems to ensure that only the filtering strength and propagation speed influence the results. The experiments were carried out on a MacBook Pro M3 with 32GB of memory. A description of the models is available in the technical appendix, and the sourcecodes are available in the supplementary material. Figure 4 shows the cumulative number of instances solved within a virtual time limit for the three problems. The appendix also presents a pairwise comparison of solvers on the number of backtracks for instances that were commonly solved.
RCPSP-CPR
@@
The Resource Constrained Project Scheduling Problem with Consumption and Production of Resources (RCPSP-CPR) is a variant of the RCPSP with additional reservoir resources that must always be kept positive (Koné et al. 2013). Each task is present and, in addition to classical renewable resource consumption (modeled with pulse functions), tasks also consume a specified amount of reservoir resources at their start and produce a specified amount at their end (modeled with stepAtStart and stepAtEnd). The objective is to minimize the makespan. We experiment on 55 instances with 15 tasks adapted from (Koné et al. 2013). A time limit of 600s is set per instance for finding and proving optimality. As shown in Figure 4, MaxiCP solves more instances than Gecode and roughly the same set of instances as CPOptimizer. A backtrack analysis between CPOptimizer and MaxiCP reveals that the number of backtracks is nearly identical, suggesting that CPOptimizer is simply faster on this problem, as both solvers achieve similar levels of filtering on this problem.
SMIC
The Single Machine with Inventory Constraints (SMIC) problem was introduced in (Davari et al. 2020). Each task has a release date, a fixed duration, and a positive or negative inventory consumption that occurs at the start of the activity (modeled with stepAtStart). The inventory starts at a given value and must stay within a specified range. Activities cannot overlap in time. The objective is to minimize the makespan. We experiment on instances with 10 tasks from (Davari et al. 2020). The non-overlap constraint was enforced using the generalized cumulative constraint in MaxiCP (modeled using pulse) and Gecode. A time limit of 600s is set per instance for finding and proving optimality. As shown in Figure 4 CPOptimizer solves more instances. A backtrack analysis showed that CPOptimizer requires significantly fewer backtracks on this problem, while Gecode and MaxiCP are not on par for this problem in terms of time and number of backtracks. Unfortunately the results for CPOptimizer for this problem are biased as it appears to reformulate the cumulative function for modeling the non-overlap into a standard disjunctive constraint for which strong dedicated filtering exist such as (Vil´ım 2004).
MESP
The Maximum Energy Scheduling Problem (MESP) involves optional tasks with variable durations and resource demands, which can be positive or negative, on a single resource with fixed capacity. The objective is to maximize the total positive energy consumption of the resources while ensuring that the resource capacity is not exceeded at any time. All filtering rules are beneficial for this problem, making it the most relevant benchmark for evaluating the proposed filtering algorithm. We generated 60 instances with a number of tasks ranging between 6 and 12800. Given the large number of tasks in some instances, this experiment is designed to test the scalability of the filtering. A time limit of 2,000s and a memory limit of 16GB is set per instance for finding a feasible solution.
As can be observed in Figure 4, our approach performs very well with more instances solved. Both Gecode and CP Optimzer fail to solve largest instances due to timeouts and memory limits respectively. A backtrack analysis shows that our approach avoids backtracking, unlike the other solvers, and that CP Optimizer requires fewer backtracks than Gecode although it solves less instances.
@@
Figure 4: Number of solved instances in function of time for the RCPSP-CPR, SMIC and MESP Problems.
Conclusion
We explained how the GeneralizedCumulative constraint can be used to implement modeling paradigm of cumulative functions for scheduling problems. We also introduced a time-tabling filtering algorithm for the GeneralizedCumulative constraint. The experimental results show that the proposed approach is scalable and competitive with other solvers on three cumulative scheduling problems with varying characteristics.

### v2

@@ 2 $p_{i}\leftarrow\mathit{true}$
Another optimization is to avoid considering some tasks when performing the timetabling algorithm. Indeed, fixed tasks that occur before the start of the earliest unfixed task or after the end of the latest unfixed task are not useful for filtering. Thus, such tasks do not need to be processed which reduces the number of time points in the profile. To do so, we use the same process as the Fruitless Fixed Tasks Removal of (GayHS15).
@@
  1. (1)
    When processing task $A$ at time point $1$, its maximum height $\overline{c}_{A}$ is adjusted to $1$ by the Height rule.
  2. (2)
    When processing task $B$ at time point $2$, the task is detected as non-feasible at this time point by the Forbid rule. Its minimum start $\underline{s}_{B}$ is thus pushed to the next time point at time $3$ where the adjustment stops as the task is feasible at this time. This also updates its minimum end to $\underline{e}_{B}=6$ due to the internal relation $\underline{e}_{B}\geq\underline{s}_{B}+\underline{d}_{B}$.
  3. (3)
    When processing task $C$ at time point $4$, the task is detected as mandatory. The following adjustments are done:
    1. (a)
      The task is set to present.
    2. (b)
      Its maximum starting time is set to $\overline{s}_{C}=4$ which also updates its maximum ending time to $\overline{e}_{C}=7$ due to the internal relation $\overline{e}_{C}\leq\overline{s}_{C}+\overline{d}_{C}$.
    3. (c)
      Its minimum ending time is set to $\underline{e}_{C}=5$.
    4. (d)
      As the value overload=1overload=-1 is negative, the maximum height of the task is set to $\overline{c}_{C}=-1$.
    Note that even if task $B$ is processed before task $C$ and its fixed part increased, task $C$ is adjusted based on the profile range which has been computed based on the state of task $B$ at the start of the propagation.
@@
AA:
: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=1,\allowbreak p=\mathit{true}\rangle$
$B$:
: $\langle s=[3,4],\allowbreak d=[3,4],\allowbreak e=[6,7],\allowbreak c=2,\allowbreak p=\mathit{true}\rangle$
$C$:
: $\langle s=[3,4],\allowbreak d=[1,3],\allowbreak e=[5,7],\allowbreak c=[-2,-1],\allowbreak p=\mathit{true}\rangle$
Figure 5 shows the state of the tasks as well as the adjustments done.
Figure 5. Representation of the tasks of Example 1 after timetabling. Adjustments are shown as gray arrows when applicable.
4. Related Work
Some forms of the generalized cumulative constraint are available in both open-source and commercial constraint programming systems. We distinguish between those that use a multi-resource API, based on the work of (BeldiceanuC02), and those that offer a conditional task interval API with cumulative functions.
4.1. Generalized multi-resource constraint
In (BeldiceanuC02), a timetabling filtering algorithm for the GeneralizedCumulative constraint with $k$ alternative resources was introduced. This constraint does not rely on conditional task intervals or cumulative functions (which were only introduced later in (LaborieRSV09)); instead, it operates directly on arrays of integer variables. The signature of the constraint is:
$cumulatives(s[1,n],d[1,n],e[1,n],c[1,n],m[1,n],\underline{C}[1,k],\overline{C}[1,k])$
where $s$, $d$, $e$, and $c$ denote respectively the start times, durations, end times, and consumptions (which can be positive or negative). The variables $m$ indicate, for each task, the resource on which it executes, with domain on $\{1,\ldots,k\}$. Finally, $\underline{C}[j],\overline{C}[j]]$ specifies the capacity range of resource $j$.
Although this constraint does not directly support conditional task intervals, it is straightforward to connect the two modeling approaches. Indeed, a dummy resource can be used to indicate if a task is not present in the multi-resource formulation. Conversely, the conditional task approach supports multiple resources by using an additional alternative constraint (LaborieR08) that represent the choice of alternative resources. This requires duplicating the activity once for each alternative resource, but it has the advantage that each alternative task can be filtered independently during propagation. In contrast, the multi-resource formulation propagation can only exploit the minimum over the different possible starts across all resources, which can lead to a weaker filtering.
The filtering algorithm of (BeldiceanuC02) proceeds by collecting a set of tasks at relevant time points before performing propagation. This algorithm is implemented in the open-source solver Gecode (schulte2006gecode) and according to its documentation, the same algorithm is also implemented in Sicstus (sicstusmanual). The algorithm presents some substantial differences with our algorithm.
First, our approach is closer to the strategy described in (GayHS15) for the classical cumulative constraint, which constructs a profile and then tests each task against it. We rely on the Profile data structure (GingrasQ16) to prune the tasks forward and backward, thus reducing the number of propagation calls required to reach a fixpoint. In contrast to (GayHS15), which represents the profile as a simple list of rectangles, our algorithm can retrieve the starting rectangle in $\mathcal{O}(1)$ thanks to the Profile. Consequently, although the worst-case quadratic time complexity per call remains similar to that of (BeldiceanuC02) and (GayHS15), our algorithm can be faster in practice. Second, our algorithm can possibly prune more the height and the duration attributes of the tasks. These filtering differences are detailed next.
Backward propagation
For the adjustment of the maximum end time ($\overline{e}$) of tasks due to the Forbid rule, when a task requires two or more consecutive adjustments, our algorithm performs all the adjustments in a single call to the timetabling algorithm compared to the version of (BeldiceanuC02) where only one adjustment is done per call. Indeed, in (BeldiceanuC02), detection and adjustment are performed only in a forward fashion. Thus, the filtering of the maximum end time of tasks is done only in regards to the current end of the task time window. This means that if a task has its maximum end time adjusted, whether or not the maximum end time can be further adjusted will be considered in the next call to the filtering algorithm. In practice, this may lead to situations where the filtering algorithm of (BeldiceanuC02) needs several consecutive calls in order to completely adjust the end time of a task and reach a fixed point. In contrast the algorithm presented in this paper does such adjustments in a single pass as illustrated in Example 1.
Example 0.
Let us consider three tasks:
  • Task A is fixed and present, starts at 4, ends at 5 and has a height of 1.
  • Task B is fixed and present, starts at 7, ends at 8 and has a height of 1.
  • Task C is present, has a minimum start of 0, a maximum end of 10, a fixed length of 3 and a height of 1.
The maximum capacity of the resource $\overline{C}$ is 1. Figure 6 shows the three tasks before propagation. The dashed blue rectangle represents the time window $\underline{s},\overline{e})$ of task C.
Figure 6. Tasks of example 1
During the backward propagation of our timetabling algorithm, task C is detected as infeasible at times 8 then 5, and its maximum end time is successively updated to 7 then 4 in a single call. The propagation from (BeldiceanuC02) iterates the profile in one direction and detects the infeasibility of task C only at time 8 and adjusts its maximum end at 7 during the first call. A second call to the algorithm is necessary in order to adjust the maximum end at time 4.
Height Adjustment Difference
For height adjustment, the difference occurs in the case where a task has no fixed part (second case of the Height filtering rule). In this case, our algorithm considers the maximum height available over the minimum overlapping interval of the task (gay2015time).
The propagation from (BeldiceanuC02) uses another height adjustment rule (see Alg. 4 in their paper). It considers the maximum available height only if the task examined has both its earliest completion time ($\underline{e}$) and its latest start time ($\overline{s}$) in the same event interval. That means that if one or more other tasks affect the available height between the $\underline{e}$ and $\overline{s}$ of a task, leading them to occur during different events, the height of the task is not adjusted. Example 2 illustrates this.
Example 0.
Let us consider three tasks:
  • Task A is fixed and present, starts at 4, ends at 12 and has a height of 2.
  • Task B is fixed and present, starts at 6, ends at 10 and has a height of -1.
  • Task C has a minimum start of 0, a maximum end of 16 and a fixed length of 6. Its height is in the interval $[1,4]$
The maximum capacity of the resource $\overline{C}$ is 4. Figure 7 shows the three tasks before propagation. The dashed blue rectangle represents task C maximum energy. The dotted black line shows the minimum profile $\underline{P}$.
Figure 7. Tasks of example 2
During the execution of the timetabling algorithm, the maximum height of task C is adjusted to 3 as this value corresponds to the maximum available height over its minimum overlapping interval (from time 5 to time 10). This filtering does not occur with the GeneralizedCumulative constraint from (BeldiceanuC02) as the $\underline{e}_{C}$ and $\overline{s}_{C}$ of task C are not in the same event due to the presence of task B. If the task B is removed, then both algorithms obtain the same filtering as the height of task C is adjusted to 2.
Length Adjustment Difference
The difference in filtering on the maximum length adjustment of the tasks occurs if the profile prevents a task to be present in at least two different parts of its time window. In this case, the algorithm presented in this paper identifies the longest interval in the profile where the task can fit and adjusts the task’s maximum length accordingly.
In contrast, the filtering of (BeldiceanuC02) one single conflict point at a same time. When encountering conflicting time point in the profile, the task maximum length is updated based on the maximum length between the conflicting time point and either the minimum start or the maximum end of the task. Example 3 illustrates this.
Example 0.
Let us consider three tasks:
  • Task A is fixed and present, starts at 3, ends at 6 and has a height of 3.
  • Task B is fixed and present, starts at 10, ends at 14 and has a height of 3.
  • Task C has a minimum start of 0, a maximum end of 16 and a height of 2. Its length is in the interval $[2,16]$
The maximum capacity of the resource $\overline{C}$ is 4. Figure 8 shows the tree tasks before propagation. The dashed blue rectangle represents task C maximum energy. The black dotted line shows the maximum capacity ($\overline{C}$). In this example, both tasks A and B prevent task C to be present at the same time.
Figure 8. Tasks of example 3
During the execution of the timetabling algorithm, the maximum length of task C is adjusted to 4 as it is the length of the longest interval where the task can be placed (from time 6 to time 10). When filtering with the propagation algorithm from (BeldiceanuC02), the length is instead adjusted to 10, since this corresponds to both the length available after task A and the length available before task B.
4.2. Closed source solvers
Both CPOptimizer (laborie2018ibm) and OptalCP (OptalCP) have an implementation of the cumulative functions modeling paradigm but the details of their filtering algorithms has not been published and both solvers are closed-source. Therefore, the exact filtering used for their range constraints on cumulative functions is difficult to assert. Based on the filtering and the number of backtracks on some examples, we observed that CPOptimizer achieves a filtering similar to the Gecode implementation when a cumulative function requires a generalized cumulative constraint. For specific cases, like when a cumulative function models a non-overlap constraint, the constraint is reformulated by CPOptimizer into a standard disjunctive constraint for which strong dedicated filtering exists (such as edge-finding, not-first/not-last, etc.).
Contrary to ours, the CPOptimizer modeling API does not allow cumulative functions with a negative height. Indeed, while cumulative tasks can have a negative contribution, a cumulative function with a negative height will cause a failure in CPOptimizer. This is equivalent to having a constraint $f\geq 0$ on any cumulative function $f$ used in the model. It is not known whether this limitation is due to a design choice or a limitation in the filtering algorithm used for the constraint.
5. Experiments
@@
Our approach and filtering algorithm are open source and available in MaxiCP (MaxiCP2024), a Java solver that extends MiniCP (MiniCP). We compare it with the closed-source state-of-the-art commercial solver CPOptimizer (laborie2018ibm) as well as the open-source solver Gecode (schulte2006gecode) that implements a version of the GeneralizedCumulative constraint of (BeldiceanuC02). As cumulative functions are not available in Gecode, we carefully modeled them in a way similar to how they would be flattened by Algorithm 1. Three cumulative problems that involve reservoirs or negative consumptions are considered. We use the same fixed static search strategy for all solvers and problems to ensure that only the filtering strength and propagation speed influence the results. The experiments were carried out on a MacBook Pro M3 with 32GB of memory. The source code of the models is available in the supplementary material.
For each problem, we report the results as a plot of the cumulative number of instances solved within a virtual time limit. We also provide a pairwise comparison of solvers on the number of backtracks for instances that were commonly solved.
Each model involves a set of conditional task intervals $i\in T$. Recall that the attributes for each interval are $s_{i}$ the start time, $d_{i}$ the duration, $e_{i}$ the end time, $p_{i}$ the status of a task (present or absent) when applicable.
5.1. The RCPSP with Consumption and Production of Resources (RCPSP-CPR)
@@
The Resource-Constrained Project Scheduling Problem with Consumption and Production of Resources (RCPSP-CPR) (kone2013comparison) is an extension of the classic Resource-Constrained Project Scheduling Problem (RCPSP) (Dike_RCPSP; DING_2023_RCPSP) with additional reservoir resources that must always be kept positive. Each task is present and, in addition to classical renewable resource consumption (modeled with pulse functions), tasks also consume a specified amount of reservoir resources at their start and produce a specified amount at their end (modeled with stepAtStart and stepAtEnd). The objective is to minimize the makespan.
Model
$T$ denotes the set of tasks, $Rn$ is the set of renewable resources, $Rs$ is the set of reservoir resources, and $P$ is the set of precedences between pairs of tasks. The capacity of each renewable resource $k\in Rn$ is denoted $C_{k}$ and the initial level of each reservoir resource $u\in Rs$ is denoted $C^{r}_{u}$. Each task $i$ consumes $c_{ki}$ unit of the renewable resource $k\in Rn$, consumes $c^{-}_{ui}$ unit of reservoir resource $u\in Rs$ at its start time and produces $c^{+}_{ui}$ unit of reservoir resource $u\in Rs$ at its end time.
The model is written as:
(11) $\mathit{minimize}\;\max\limits_{i\in T}e_{i}$
subject to
(12) $e_{i}\leq s_{j}$ $\forall(i,j)\in P$
(13) $ren_{k}=\sum\limits_{i\in T}pulse(i,c_{ki})$ $\forall k\in Rn$
[…]
(15) $ren_{k}\leq C_{k}$ $\forall k\in Rn$
(16) $res_{u}\geq 0$ $\forall u\in Rs$
The objective (11) is to minimize the makespan. The constraints (12) ensure the precedences between tasks. The renewable and reservoir resources are constrained in (13) and (14) respectively. The resource constraints are posted at (15) and (16) respectively.
The search consists in a static binary branching that selects variables in the order in which tasks are declared in the instance file. The first un-assigned start variables is selected and assigned to its minimum value in the left branch. The right branch removes the minimum value from the domain.
Results
We experiment on 55 instances with 15 tasks adapted from (kone2013comparison). A time limit of 600s is set per instance for finding and proving optimality. As shown in Figure 9, MaxiCP solves more instances than Gecode and roughly the same set of instances as CPOptimizer.
@@ Results
Figure 9. Number of solved instances in function of time for the RCPSP-CPR Problem.
Figure 10 reports for each pair of solvers, for each instance solved commonly by both solvers, an $x,y$- plot of the number of backtracks. The comparison between CPOptimizer and MaxiCP reveals that the number of backtracks is nearly identical, suggesting that CPOptimizer is simply faster on this problem, as both solvers achieve similar levels of filtering on this problem, but MaxiCP requires less backtracks than Gecode form some of the instances.
Figure 10. Failure comparison for the RCPSP-CPR
5.2. The Single Machine with Inventory Constraints (SMIC)
The Single Machine with Inventory Constraints (SMIC) problem was introduced in (smic). Each task has a release date $release_{i}$, a fixed duration, and a positive or negative inventory consumption $invent_{i}$ that occurs at the start of the activity (modeled with stepAtStart). The inventory starts at a given value $\mathit{initInventory}$ and must stay within a specified range ($0,\mathit{capaInventory}]$). Activities cannot overlap in time. The objective is to minimize the makespan.
Model
The model is written as:
(17) $\mathit{minimize}\;\max_{i\in T}e_{i}$
subject to
(18) $s_{i}\geq\mathit{release_{i}},$ $\forall i\in T$
(19) $res=\mathit{step}(0,\mathit{initInventory})+\sum_{i\in T}\mathit{stepAtStart}(i,invent_{i}),$ $\forall i\in T$
(20) $0\leq res\leq\mathit{capaInventory}$
(21) $\mathit{noOverlap}=\sum_{i\in T}pulse(i,1),$ $\forall i\in T$
(22) $\mathit{noOverlap}\leq 1$
The objective (17) is to minimize the makespan (i.e., the time at which all tasks are completed). The release dates are ensured by the relation (18). The reservoir resource is constrained in relations (19) and (20) to ensure that the capacity of the inventory is not violated. The noOverlap constraint is enforced with the relations (21) and (22) to ensure the no-overlapping of activities. Notice that those are modeled using cumulative functions rather than with global dedicated non-overlap constraints on task intervals to ensure that the results are solely impacted by the filtering of the cumulative constraint rather than by the possibly different filtering of the non-overlap constraint. This, unfortunately, did not appear possible for CPOptimizer which recognizes that the cumulative function is used to model a non-overlap constraint and automatically reformulates it using a non-overlap constraint with stronger filtering.
The search consists in a static branching heuristic that selects variables in the order in which tasks are declared in the instance file. The first un-assigned start variables is selected and assigned to its minimum value in the left branch. The right branch removes the minimum value from the domain.
Results
We experiment on instances with 10 tasks from (smic). A time limit of 600s is set per instance for finding and proving optimality. As shown in Figure 11 CPOptimizer solves more instances while MaxiCP and Gecode roughly have the same behavior on this problem.
Figure 11. Number of solved instances in function of time for the SMIC Problem
Figure 12 reports for each pair of solvers, for each instance solved commonly by both solvers, an $x,y$- plot of the number of backtracks. This analysis shows that CPOptimizer requires significantly fewer backtracks on this problem, while Gecode and MaxiCP are not on par for this problem in terms of time and number of backtracks. Unfortunately the results for CPOptimizer for this problem are biased as it appears to reformulate the cumulative function for modeling the non-overlap into a standard disjunctive constraint for which strong dedicated filtering exist such as (vilim2004nlog).
Figure 12. Failure comparison for the SMIC
5.3. The Maximum Energy Scheduling Problem (MESP)
The Maximum Energy Scheduling Problem (MESP) consists in scheduling optional tasks with variable durations and resource demands, which can be positive or negative, on a single resource with fixed capacity $\overline{C}$. The energy of a task is the product of the duration and the demand. The objective is to maximize the total positive energy consumption of the resources while ensuring that the resource capacity is not exceeded at any time. All filtering rules are beneficial for this problem, making it the most relevant benchmark for evaluating the proposed filtering algorithm.
The model is written as:
(23) $\mathit{maximize}\;\sum\limits_{i\in T\mid p_{i}=\mathit{true}}\max(c_{i},0)\cdot d_{i}$
subject to
(24) $\mathit{resource}=\sum\limits_{i\in T}\mathit{pulse}(i,c_{i})$
(25) $\mathit{resource}\leq\overline{C}$
The objective (23) is to maximize the sum of the positive energies of the tasks that are present. The constraint (24) sets up the reservoir resource and its capacity is constrained by (25).
Note that the resource demand variables of tasks $c_{i}$ may have negative values in their domain in this problem. As explained in Filtering Differences section, CPOptimizer does not allow cumulative functions to have a negative height. In order to accommodate this restriction, the model for CPOptimizer has been adapted in the following way: The cumulative function resourceresource is shifted up in height at the time 0 by a step function of height $neg^{*}=\sum_{i\in T}-\min(\underline{c}_{i},0)$ which corresponds to the sum of all maximum negative values of the tasks resources demands. The capacity $\overline{C}$ of the resource is also shifted by the same value. In formal terms, constraint (24) is changed to: $resource=step(0,neg^{*})+\sum\limits_{i\in T}pulse(i,c_{i})$ and constraint (25) is changed to $resource\leq\overline{C}+neg^{*}$. This ensures that the cumul function resourceresource remains non negative, as required by CPOptimizer.
The search heuristic consists in selecting tasks in the fixed order of their declaration in the instance file. For each task, the following variables are fixed in this order:
  • Its presence $p_{i}$ to true
  • Its demand $c_{i}$ to the maximum value in the domain
  • Its duration $d_{i}$ to the maximum value in the domain
  • Its end time $e_{i}$ to the maximum value in the domain
Two branches are created: The left branch fixes the value while the right branch removes it from the domain. The search stops at the first solution found.
Results
We generated 60 instances with a number of tasks ranging between 6 and 12800. A time limit of 2,000s and a memory limit of 16GB is set per instance for finding a feasible solution. Given the large number of tasks in some instances, this experiment is designed to test the scalability of the filtering. As can be observed in Figure 13, our approach performs very well with more instances solved. Both Gecode and CP Optimzer fail to solve largest instances due to timeouts and memory limits respectively.
Figure 13. Number of solved instances in function of time for the MESP Problem.
Figure 14 reports for each pair of solvers, for each instance solved commonly by both solvers, an $x,y$- plot of the number of backtracks. As can be seen, our approach avoids backtracking, unlike the other solvers. Interestingly, CPOptimizer requires fewer backtracks than Gecode although it solves less instances.
Figure 14. Failure comparison for the MESP
5.4. Why not compare with solvers using Minizinc?
The MiniZinc modeling language provides a cumulatives constraint that allows for negative consumption. Unfortunately, with the exception of Sicstus (CarlssonM12) and Gecode (schulte2006gecode), this constraint is decomposed in other solvers. Also the MiniZinc API for this constraint does not support explicit access to the end variables. Being able to access the end variables directly rather than relying on the relation start + duration = end is crucial for the problems considered, as the end variables can be constrained independently, as explained in the next example.
Example 0.
Assume a task with the following properties: $s=[3,7]$, $d=[1,5]$, $e=[8,8]$. The mandatory part of this task is $\overline{s},\underline{e}-1]=[7,7]$. If the end variable is not available, the mandatory part inferred from $s$ and $d$ is $\overline{s},\underline{s}+\underline{d}-1]=\phi$.
As a result, the propagation when using MiniZinc for these problems can be much weaker, making the comparison unfair. This was confirmed experimentally and can be verified with the provided MiniZinc models in the appendix.
6. Conclusion
We presented an implementation of the cumulative functions modeling paradigm for scheduling problems. Our implementation applies a flattening procedure to cumulative function expressions, producing a set of activities with positive or negative resource consumption, which are then passed to a GeneralizedCumulative constraint.
We introduced a novel time-tabling filtering algorithm for the GeneralizedCumulative constraint that operates on conditional task intervals and supports negative resource consumption. This algorithm is simpler and achieves stronger pruning than the one proposed by (BeldiceanuC02).
Experimental results demonstrate that the proposed approach is scalable and competitive with other solvers on three cumulative scheduling problems with diverse characteristics, including optional tasks and negative resource consumption.
As future work, we plan to extend the filtering algorithm to generalize the Overload/Underload checking and edge-finding rules in the context of producers and consumers. We also intend to explore explanations for the generalized cumulative constraint, similar to those proposed in (schutt2011explaining), and to investigate whether a CP-SAT solver such as OR-Tools (perron_et_al) could benefit from them.